OpenStudy (anonymous):

Which statement is true for the equation 3n − 7 = 3n − 6? A. It has no solution B. It has one solution C. It has two solutions D. It has infinitely many solutions

OpenStudy (anonymous):

m=m <---means there are infinte solutions m=x <---No solution x=1 <---one solution

OpenStudy (anonymous):

Subtract 3n from both sides. \[\huge~\rm~3n−7−3n=3n−6−3n\]−7=−6 Add 7 to both sides. \[\huge~\rm−7+7=−6+7\] \[\huge~0=1\] so we have m=x case

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

so its one solution ? @Moe95

OpenStudy (anonymous):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Moe95 m=m <---means there are infinte solutions m=x <---No solution x=1 <---one solution \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (anonymous):

im lost @Moe95

OpenStudy (anonymous):

a number that doesnt equal the same number means that there is no solution 0=1 would be no solution